# Orbital orders under magnetic fields in cubic PrIr2Zn20

### v2

@@ IV.2 Phase diagram
Now, we discuss $T$$H$ phase diagram ($H\equiv|\bm{H}|$) taking into account multiple-$\bm{q}$ orders in addition to the single-$\bm{q}$ ones. We numerically minimize the free energy $F_{\rm L}+F_{\Gamma}+\delta F_{3}+\delta F_{4}+\delta F$, where $\delta F$ is due to the integration of the X point fields as discussed in Appendix B and the others are given in Eqs. (6), (10), (13), and (15). Figures 6 and 8 show phase diagrams for $\bm{H}\parallel[001]$, i.e., $\tilde{\bm{h}}=\tilde{h}_{u}(1,0)$ $(\tilde{h}_{u}\geq 0)$, and double-$\bm{q}$ configurations in the quadrupole space and real space. Figure 6(a) shows the phase diagram for $J_{1}=1$, $c=0.5$, and $b=0.1$. The exchange coupling constants are the same as used in Fig. 3: $(J_{2},J_{3},J_{4})=(0.0,0.0,1.0)$. Except for the low-field region, the phase diagram is similar to those shown in Fig. 3, showing the single-$\bm{q}$ orders I, II, and III. An orthogonal double-$\bm{q}$ configuration I2q with $\bm{\phi}_{{\bm{k}}_{1}}=(\phi_{1},0)$ and $\bm{\phi}_{{\bm{k}}_{2}}=(0,\phi_{2})$ is stabilized for finite magnetic fields at high $T$ region below the phase I. Here, we take a domain with the ordering vectors ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$. The configuration of I2q state in the quadrupole space is shown in Fig. 8(a) and its real-space one is illustrated in Fig. 8(b). At ${\bm{H}}=0$, the orthogonal double-$\bm{q}$ states with $\phi_{1}=\phi_{2}$ discussed in Sec. IV.1 are realized and degenerate in the angle $\theta=0$. As ${\bm{H}}$ increases, $\phi_{1}$ becomes larger than $\phi_{2}$. Since the phase I is the single-$\bm{q}$ order with $v_{\rm L}=0$, the symmetry breaking from the phase I to the phase I2q is caused by the emergence of $\bm{\phi}_{2}=(0,\phi_{2})$ through the second-order transition. In this sense, $\phi_{1}>\phi_{2}$ is quite natural. In the free energy, this is due to the $\delta F_{3}$, which is minimized when the L point field is parallel to $\bm{\phi}_{0}=(u_{0},0)$, while maximized when it is perpendicular. As for the phase boundary between the phase I2q and the phase II, it is of first-order, while the other phase boundaries are second-order.
At lower $T$, the induced $\Gamma$ point moment $\bm{\phi}_{0}$ favors the configuration III2q shown in Figs. 8(c) and 8(f), where $\bm{\phi}_{1}=\tfrac{1}{\sqrt{2}}\phi_{1}(1,1)$ and $\bm{\phi}_{2}=\tfrac{1}{\sqrt{2}}\phi_{2}(-1,1)$ with $\phi_{1}=\phi_{2}$. This configuration is stabilized by the nontrivial coupling discussed in the last part of Sec. IV.1. See also the last term in Eq. (39). In Fig. 6(a), an intermediate phase II2q toward the phase III2q appears at low temperatures. The quadrupole configurations in the phase II2q are “rotated” from those in the phase I2q with keeping $\bm{\phi}_{1}\perp\bm{\phi}_{2}$ as shown in Fig. 8(b). For smaller $b=0.05$ [Fig. 6(b)], the phase I, II, and the double-$\bm{q}$ orders shrink, while the phase III expands. The overall phase diagram, however, is unchanged.
As an example for more general cases, we show the phase diagram for $(b,J_{2},J_{3},J_{4})=(0.03,0.5,0.3,0.4)$ in Fig. 6(c). For this parameter set, the second-order coefficient $a_{\rm X}$ of the X point fields almost vanish at $T=0$ as shown in Fig. 1(c), the renormalization effects due to the X point fields grows rapidly at low temperatures. This causes the strong stabilization of the phase III2q in Fig. 6(c), showing reentrant behavior as a function of $H$. Interestingly, there is a thin region of the phase II2q. The phase boundary between the phase II2q and II is of the second order. Similarly, that between the small region of phase I2q and III is also second order. We note that the boundaries indicated by the dashed lines are not symmetry breaking phase boundary since the symmetry of the three double-$\bm{q}$ phases are the same as is evident from the real space configurations in Figs. 8(d)–8(f). See also the discussion in Appendix C.
For $\bm{H}\parallel[110]$, a phase I2q appears at low field region as shown in Fig. 7. In contrast to the phase I2q for $\bm{H}\parallel[001]$, the quadrupole configuration is $\bm{\phi}_{1}=(0,\phi_{1})$ and $\bm{\phi}_{2}=(\phi_{2},0)$ with $\phi_{1}>\phi_{2}$. This is natural since the neighboring single-$\bm{q}$ phase is the phase IV. The sharp increase of the critical field between I2q and IV phases are due to the fact that $a_{\rm X}$ is close to 0 in (c).
@@
Figure 8: Schematic orthogonal double-$\bm{q}$ order parameter configurations for (a) I2q: $(\theta_{1},\theta_{2})=(0,\pi/2)$ with $\phi_{1}>\phi_{2}$, (b) II2q: $(\theta_{1},\theta_{2})=(\theta,\theta+\pi/2)$ with $\phi_{1}>\phi_{2}$, and (c) III2q: $(\theta_{1},\theta_{2})=(\pi/4,3\pi/4)$ and $\phi_{1}=\phi_{2}$. Arrows indicate the order parameters $\bm{\phi}_{1}$ (L1), $\bm{\phi}_{2}$ (L2), and $\bm{\phi}_{0}$ $(\Gamma)$. Four small circles represent quadrupole moments at the A sublattice real-space positions $\bm{r}=(0,0,0),(1,0,0),(\tfrac{1}{2},\frac{1}{2},0)$, and $(\tfrac{1}{2},\frac{1}{2},1)$. Real space L point quadrupole configurations for (d) $(\theta_{1},\theta_{2})=(0,\pi/2)$ and $\phi_{1}=\sqrt{3}\phi_{2}$ in the phase I2q, (e) $(\theta_{1},\theta_{2})=(\pi/24,13\pi/24)$ and $\phi_{1}=1.2\phi_{2}$ in the phase II2q, and (f) $(\theta_{1},\theta_{2})=(\pi/4,3\pi/4)$, and $\phi_{1}=\phi_{2}$ in the phase III2q. Note that only L point quadrupole components without those at the $\Gamma$ point are shown. Skyblue and gray arrows represent the directions of ${\bm{k}}_{1}$ and $-{\bm{k}}_{2}$. For guide to eyes, some of the nearest-neighbor bonds are connected by white bonds. Circles are the A sublattice atoms in the diamond structure and their colors represent the sublattice (red, green, blue, white). For clarity, the B sublattice atoms are not drawn. Colored bonds running [01¯$1]$ 1 ] directions connects the A sublattice sites with identical quadrupole moments.
@@ V.1 Comparison with the experimental data of PrIr2Zn20
PrIr2Zn20 shows several phase transitions under magnetic field at low temperatures as reported in the early study.Onimaru et al. (2011) However, the order parameters have not been identified except for the $x^{2}-y^{2}$ type quadrupole order with the ordering vector at the L point for $\bm{H}\parallel[110]$ by the neutron scattering experimentIwasa et al. (2017). The analysis of this study is fully based on this fact. Namely, we have assumed that the leading order parameter is the quadrupole moment at the L points. Constructing the symmetry-allowed free energy consisting of them and the related degrees of freedom, we have analyzed the phase diagram for $\bm{H}\parallel[001]$ in addition to [110]. Our analysis demonstrates that a $2z^{2}-x^{2}-y^{2}$ configuration with the ordering vector at the L point is stable at high-$T$ and up to intermediate magnetic field strength for $\bm{H}\parallel[001]$. This suggests that the $2z^{2}-x^{2}-y^{2}$ order can be a promising candidate for the high-$T$ phase found for $\bm{H}\parallel[001]$ in the experimental worksOnimaru et al. (2011); Kittaka et al. (2024). As for the order parameter at zero field, the single-$\bm{q}$ configuration is not stable while the double-$\bm{q}$ I2q is realized with $\phi_{1}=\phi_{2}$. For finite ${\bm{H}}\parallel[001]$, this double-$\bm{q}$ configuration is modified and $\bm{\phi}_{1}=(\phi_{1},0)$ and $\bm{\phi}_{2}=(0,\phi_{2})$ with $\phi_{1}>\phi_{2}$ is favored. $\phi_{2}$ increases as lowering $T$ and there appear phase II2q and III2q by simultaneously rotating $\bm{\phi}_{1}$ and $\bm{\phi}_{2}$. As the cubic anisotropy $b$ increases, the influence of the renormalized fourth-order coefficients becomes larger and favors the phase II2q and III2q.
@@ V.2 Future perspectives
We have shown that single- and double-$\bm{q}$ quadrupole orders are realized in our study as a phenomenological model for PrIr2Zn20. The orthogonal double-$\bm{q}$ orders are favored by the effective fourth-order terms due to the X point fields, while the quadruple ones are not. Although we are interested in the orbital orders at the L points in this paper, quadruple-$\bm{q}$ orders at incommensurate ordering vectors along [111] directions has been studied in cubic magnets, showing hedgehog type topological magnetic structure.Okumura et al. (2020, 2022) In principle, such exotic phases are also possible for the orbital model discussed in this paper when the exchange coupling favors incommensurate ordering vectors by introducing interaction frustration or anisotropic interaction ignored in this studyHattori et al. (2024).
@@ Appendix A Multiple-q configuration
In this Appendix, we show the real-space modulations for the double-$\bm{q}$ and quadruple-$\bm{q}$ orders since they are rather complicated. The real-space configurations for the double- and quadrupole-$\bm{q}$ orders are given by Eq. (25). Here, we concentrate on the sublattice independent orders $\bm{\phi}(\bm{r})$ with $\Phi^{\alpha}=\Phi$ and $\theta^{\alpha}_{n}=\theta_{n}$ since only such orders appear in this paper. Thus, one rewrites Eq. (25) as
$\bm{\phi}(\bm{r})=\mathbb{E}_{\mathcal{N}}(\bm{r})\cdot\begin{pmatrix}\bm{\phi}_{1}\\ \bm{\phi}_{2}\\ \cdots\end{pmatrix},\ \bm{\phi}_{n}=\frac{\Phi_{n}}{\sqrt{\mathcal{N}}}\begin{pmatrix}\cos\theta_{n}\\ \sin\theta_{n}\end{pmatrix},$ (29)
@@
Here, we have assumed, without loss of generality, double-$\bm{q}$ orders with ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$. Since $\bm{r}$ is the position in the fcc lattice, there are four choices for $\mathbb{E}_{2}(\bm{r}):(\pm 1,\pm 1)$, and eight choices for $\mathbb{E}_{4}(\bm{r}):\pm(1,1,1,1),\ \pm(1,1,-1,-1),\ \pm(1,-1,1,-1)$, and ±$\pm(-1,1,1,-1)$. In Fig. 9, $\mathbb{E}_{\mathcal{N}}(\bm{r})$’s are illustrated and the colors distinguish the vectors $\mathbb{E}_{2,4}$. Note that only the A-sublattice positions in the diamond structure are shown. The unit cell for the double-$\bm{q}$ orders contains four A sublattice sites, each of which forms ferroic chain along the direction parallel to ${\bm{k}}_{3}+{\bm{k}}_{4}\parallel(0,-1,1)$ as shown in Fig. 9(a). The primitive translation vector along this direction is unchanged, while those perpendicular to it are modified to $\frac{1}{2}(2,1,1)$ and $\frac{1}{2}(2,-1,-1)$. For quadruple-$\bm{q}$ orders, the primitive translational vectors are as twice large as those for the original fcc lattice: (1,1,0), (1,0,1), and (0,1,1). The unit cell contains eight sites for the quadruple-$\bm{q}$ orders and sixteen when the B-sublattice is included.
@@
Since several double-$\bm{q}$ configurations are shown in Fig. 8, we here examine the real space configuration of the symmetric quadruple-$\bm{q}$ orders discussed in Sec. IV.1, although they are not stabilized in the $T$$H$ phase diagram in Fig. 6. It is evident that the configurations with “collinear” types such as $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,0,0)$ and $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,0,\pi)$ are partially ordered states. For example, the quadrupole moments vanish at the sites where $\sum_{n}e^{i{\bm{k}}_{n}\cdot\bm{r}}=0$ for the former as shown in Fig. 9(c). As an example of “non-collinear” case, Fig. 9(d) illustrates the configuration for $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,\pi/2,3\pi/2)$.

### v3

@@ IV.2 Phase diagram
Now, we discuss $T$$H$ phase diagram ($H\equiv|\bm{H}|$) taking into account multiple-$\bm{q}$ orders in addition to the single-$\bm{q}$ ones. We numerically minimize the free energy $F_{\rm L}+F_{\Gamma}+\delta F_{3}+\delta F_{4}+\delta F$, where $\delta F$ is due to the integration of the X point fields as discussed in Appendix B and the others are given in Eqs. (6), (10), (13), and (15). For simplify the analysis, we assume ${\boldsymbol{\phi}}_{n}^{B}={\boldsymbol{\phi}}_{n}^{A}(-1)^{p_{n}}$ valid for $J_{\rm L}<0$ and do not take into account quadruple-$\bm{q}$ orders, since their energy is much higher than single- and double-$\bm{q}$ states for the conventional fourth-order terms [Eq. (8)] as shown in Fig. 6. Figures 6 and 7 show phase diagrams for $\bm{H}\parallel[001]$, i.e., $\tilde{\bm{h}}=\tilde{h}_{u}(1,0)$ $(\tilde{h}_{u}\geq 0)$, and those for $\bm{H}\parallel[110]$ with $\tilde{\bm{h}}=\tilde{h}_{u}(1,0)$ $(\tilde{h}_{u}\leq 0)$, respectively. The representative double-$\bm{q}$ configurations in the quadrupole space and real space are also illustrated in Fig. 8.
@@
Figure 6(a) shows the phase diagram for $J_{1}=1$, $c=0.5$, and $b=0.1$. The exchange coupling constants are the same as used in Fig. 3: $(J_{2},J_{3},J_{4})=(0,0,1)$. Except for the low-field region, the phase diagram is similar to those shown in Fig. 3, showing the single-$\bm{q}$ orders I, II, and III. An orthogonal double-$\bm{q}$ configuration I2q with $\bm{\phi}_{{\bm{k}}_{1}}=(\phi_{1},0)$ and $\bm{\phi}_{{\bm{k}}_{2}}=(0,\phi_{2})$ is stabilized for finite magnetic fields at high $T$ region below the phase I. Here, we take a domain with the ordering vectors ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$. The configuration of I2q state in the quadrupole space is shown in Fig. 8(a) and its real-space one is illustrated in Fig. 8(b). At ${\bm{H}}=0$, the orthogonal double-$\bm{q}$ states with $\phi_{1}=\phi_{2}$ discussed in Sec. IV.1 are realized. As $H$ increases, $\phi_{1}$ becomes larger than $\phi_{2}$. Since the phase I is the single-$\bm{q}$ order with $v_{\rm L}=0$, the symmetry breaking from the phase I to the phase I2q is caused by the emergence of $\bm{\phi}_{2}=(0,\phi_{2})$ through the second-order transition. In this sense, $\phi_{1}>\phi_{2}$ is quite natural. In the free energy, this is due to the $\delta F_{3}$, which is minimized when the L point field is parallel to $\bm{\phi}_{0}=(u_{0},0)$, while maximized when it is perpendicular. As for the phase boundary between the phase I2q and the phase II, it is of first-order, while the other phase boundaries are second-order.
At lower $T$, the induced $\Gamma$ point moment $\bm{\phi}_{0}$ favors the configuration III2q shown in Figs. 8(c) and 8(f), where $\bm{\phi}_{1}=\tfrac{1}{\sqrt{2}}\phi_{1}(1,1)$ and $\bm{\phi}_{2}=\tfrac{1}{\sqrt{2}}\phi_{2}(-1,1)$ with $\phi_{1}=\phi_{2}$. This configuration is stabilized by the fifth-order coupling discussed in the last part of Sec. IV.1. See also the last term in Eq. (39). In Fig. 6(a), an intermediate phase II2q toward the phase III2q appears at low temperatures. The quadrupole configurations in the phase II2q are “rotated” from those in the phase I2q with keeping $\bm{\phi}_{1}\perp\bm{\phi}_{2}$ as shown in Fig. 8(b). For smaller $b=0.05$ [Fig. 6(b)], the phase I, II, and the double-$\bm{q}$ orders shrink, while the phase III expands. The overall phase diagram, however, is unchanged.
@@
Figure 8: Schematic orthogonal double-$\bm{q}$ order parameter configurations for (a) I2q: $(\theta_{1},\theta_{2})=(0,\pi/2)$ with $\phi_{1}>\phi_{2}$, (b) II2q: $(\theta_{1},\theta_{2})=(\theta,\theta+\pi/2)$ with $\phi_{1}>\phi_{2}$, and (c) III2q: $(\theta_{1},\theta_{2})=(\pi/4,3\pi/4)$ and $\phi_{1}=\phi_{2}$. Arrows indicate the order parameters $\bm{\phi}_{1}$ (L1), $\bm{\phi}_{2}$ (L2), and $\bm{\phi}_{0}$ $(\Gamma)$. Four small circles represent the quadrupole moments at the $A$ sublattice real-space positions $\bm{r}=(0,0,0),(1,0,0),(\tfrac{1}{2},\frac{1}{2},0)$, and $(\tfrac{1}{2},\frac{1}{2},1)$. Real space L point quadrupole configurations for (d) $(\theta_{1},\theta_{2})=(0,\pi/2)$ and $\phi_{1}=\sqrt{3}\phi_{2}$ in the phase I2q, (e) $(\theta_{1},\theta_{2})=(\pi/24,13\pi/24)$ and $\phi_{1}=1.2\phi_{2}$ in the phase II2q, and (f) $(\theta_{1},\theta_{2})=(\pi/4,3\pi/4)$, and $\phi_{1}=\phi_{2}$ in the phase III2q. Note that only L point quadrupole components without those at the $\Gamma$ point are shown. For guide to eyes, some of the nearest-neighbor bonds are connected by white bonds. Circles are the $A$ sublattice atoms and their colors represent the sublattice (red, green, blue, white). For clarity, the $B$ sublattice atoms are not drawn. Colored bonds running [01¯$1]$ 1 ] directions connects the $A$ sublattice sites with the same quadrupole moments.
As an example of more general cases, we show the phase diagram for $(b,J_{2},J_{3},J_{4})=(0.03,0.5,0.3,0.4)$ in Fig. 6(c). For this parameter set, the second-order coefficient $a_{\rm X}$ of the X point fields almost vanish at $T=0$ as shown in Fig. 1(c), the renormalization effects due to the X point fields grows rapidly at low temperatures. This causes the strong stabilization of the phase III2q in Fig. 6(c), showing reentrant behavior as a function of $H$. Note, however, the Landau theory is not applicable to such low temperature region far away from $T\sim T_{c}$. There is a thin region of the phase II2q. The phase boundary between the phase II2q and II is of the second order. Similarly, that between the small region of phase I2q and III is also second order. We note that the boundaries indicated by the dashed lines are not symmetry breaking phase boundary since the symmetry of the three double-$\bm{q}$ phases are the same as is evident from the real space configurations in Figs. 8(d)–8(f). See also the discussion in Appendix C.
For $\bm{H}\parallel[110]$, a phase I2q appears at low-field region as shown in Fig. 7. In contrast to the phase I2q for $\bm{H}\parallel[001]$, the quadrupole configuration is $\bm{\phi}_{1}=(0,\phi_{1})$ and $\bm{\phi}_{2}=(\phi_{2},0)$ with $\phi_{1}>\phi_{2}$. This is natural since the neighboring single-$\bm{q}$ phase is the phase IV. The sharp increase of the critical field between I2q and IV phases in Fig. 7(c) is due to the fact that $a_{\rm X}$ is close to zero.
@@ V.1 Comparison with the experimental data of PrIr2Zn20
PrIr2Zn20 shows several phase transitions under magnetic field at low temperatures as reported in the early study.Onimaru et al. (2011) However, the order parameters have not been identified except for the $x^{2}-y^{2}$ type quadrupole order with the ordering vector at the L point for $\bm{H}\parallel[110]$ by the neutron scattering experimentIwasa et al. (2017). The analysis of this study is based on this experimental fact. Namely, we have assumed that the leading order parameter is the quadrupole moment at the L points. Constructing the symmetry-allowed free energy consisting of them and the related degrees of freedom, we have analyzed the phase diagram for $\bm{H}\parallel[001]$ in addition to that for $\bm{H}\parallel[110]$. Our analysis demonstrates that a $2z^{2}-x^{2}-y^{2}$ configuration with the ordering vector at the L point is stable at high-$T$ and up to intermediate magnetic field strength for $\bm{H}\parallel[001]$. This suggests that the $2z^{2}-x^{2}-y^{2}$ order can be a promising candidate for the high-$T$ phase found for $\bm{H}\parallel[001]$ in the experimental worksOnimaru et al. (2011); Kittaka et al. (2024). As for the order parameter at zero field, the single-$\bm{q}$ configuration is not stable while the double-$\bm{q}$ I2q is realized with $\phi_{1}=\phi_{2}$. For finite ${\bm{H}}\parallel[001]$, this double-$\bm{q}$ configuration is modified and $\bm{\phi}_{1}=(\phi_{1},0)$ and $\bm{\phi}_{2}=(0,\phi_{2})$ with $\phi_{1}>\phi_{2}$ is favored. $\phi_{2}$ increases as lowering $T$ and there appear phase II2q and III2q by simultaneously rotating $\bm{\phi}_{1}$ and $\bm{\phi}_{2}$. As the cubic anisotropy $b$ increases, the influence of the renormalized fourth-order coefficients becomes larger and favors the phase II2q and III2q.
@@ V.2 Future perspectives
We have shown that single- and double-$\bm{q}$ quadrupole orders are realized in our study as a phenomenological model for PrIr2Zn20. The orthogonal double-$\bm{q}$ orders are favored by the effective fourth-order terms arising from the renormalization due to the X point fields, while the quadruple ones are not. Although we are interested in the orbital orders at the L points in this paper, quadruple-$\bm{q}$ orders at incommensurate ordering vectors along [111] directions has been studied in cubic magnets, showing hedgehog type topological magnetic structure.Okumura et al. (2020, 2022) In principle, such exotic phases are also possible for the orbital model discussed in this paper when the exchange coupling favors incommensurate ordering vectors by introducing interaction frustration or anisotropic interaction ignored in this studyHattori et al. (2024). In this sense, it is interesting to explore possibilities of incommensurate orbital orders as the candidates for the high-$T$ and intermediate field phase instead of the phase I proposed in this paper as studied in magnetic ordersD. Bergman and Balents (2007); S. Gao and Rüegg (2017); Rosales et al. (2022).
@@ Appendix A Multiple-q configuration
In this Appendix, we show the real-space modulations for the double-$\bm{q}$ and quadruple-$\bm{q}$ orders since they are rather complicated. The real-space configurations for the double- and quadrupole-$\bm{q}$ orders are given by Eq. (25). Here, we concentrate on the $A$ sublattice moments, since those for the $B$ sublattice are obtained by ${\boldsymbol{\phi}}_{n}\to{\boldsymbol{\phi}}_{n}(-1)^{p_{n}}$. Thus, one can rewrite Eq. (25) as
$\bm{\phi}^{A}(\bm{r})=\mathbb{E}_{\mathcal{N}}(\bm{r})\cdot\begin{pmatrix}\bm{\phi}_{1}\\ \bm{\phi}_{2}\\ \cdots\end{pmatrix},\ \bm{\phi}_{n}=\frac{\Phi_{n}}{\sqrt{\mathcal{N}}}\begin{pmatrix}\cos\theta_{n}\\ \sin\theta_{n}\end{pmatrix},$ (29)
@@
Here, we have assumed, without loss of generality, double-$\bm{q}$ orders with ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$. Since $\bm{r}$ is the position in the fcc lattice, there are four choices for $\mathbb{E}_{2}(\bm{r}):(\pm 1,\pm 1)$, and eight choices for $\mathbb{E}_{4}(\bm{r}):\pm(1,1,1,1),\ \pm(1,1,-1,-1),\ \pm(1,-1,1,-1)$, and ±$\pm(-1,1,1,-1)$. In Fig. 9, $\mathbb{E}_{\mathcal{N}}(\bm{r})$’s are illustrated and the colors distinguish the vectors $\mathbb{E}_{2,4}$. Note that only the $A$-sublattice positions in the diamond structure are shown. The unit cell for the double-$\bm{q}$ orders contains four $A$ sublattice sites, each of which forms ferroic chain along the direction parallel to ${\bm{k}}_{3}+{\bm{k}}_{4}\parallel(0,-1,1)$ as shown in Fig. 9(a). The primitive translation vector along this direction is unchanged, while those perpendicular to it are modified to $\frac{1}{2}(2,1,1)$ and $\frac{1}{2}(2,-1,-1)$. For quadruple-$\bm{q}$ orders, the primitive translational vectors are as twice large as those for the original fcc lattice: (1,1,0), (1,0,1), and (0,1,1). The unit cell contains eight sites for the quadruple-$\bm{q}$ orders and sixteen when the $B$-sublattice sites are included.
@@
Since several double-$\bm{q}$ configurations are shown in Fig. 8, we here examine the real space configuration of the symmetric quadruple-$\bm{q}$ orders discussed in Sec. IV.1, although they are not stabilized in the $T$$H$ phase diagram in this paper. It is evident that the configurations with “collinear” types such as $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,0,0)$ and $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,0,\pi)$ are partially ordered states. For example, the quadrupole moments vanish at the sites where $\sum_{n}e^{i{\bm{k}}_{n}\cdot\bm{r}}=0$ for the former as shown in Fig. 9(c). As an example of “non-collinear” case, Fig. 9(d) illustrates the configuration for $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,\pi/2,3\pi/2)$.